Grade 11Physical Chemistry

what is the volume occupied by one mole of hcl at stp.

SATYADEB HOTA
8 Years agoGrade 11

1 Answer

Surjeet Singh Kushwah
8 Years ago
PV=nRT
10^5*V=1*8.314*273. (at STP pressure = 10^5)
V~44l
Dear student for any don't hesitate to ask
Thank you